OpenStudy (anonymous):

A delivery truck made 8 stops in one neighborhood. By stop 5, there were 11 packages delivered. How many packages were delivered by stop 3? A. 6 B. 5 C. 3 D. 1

OpenStudy (anonymous):

help someone i will give u an medal

OpenStudy (anonymous):

A

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

thxs tho

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

was i right or do u not know yet all u do is just look at the graph

OpenStudy (anonymous):

@israa88 how did you get 6?

OpenStudy (anonymous):

wait its D

OpenStudy (anonymous):

i got 5

OpenStudy (anonymous):

how ? u might be right i said i think its a from the beginning so i am not 100% sure

OpenStudy (anonymous):

well you must be right you got best response soo

OpenStudy (anonymous):

but any who i did 8-5=3*2=6 11packages - 6=5 but i did it wrong , i guess

OpenStudy (anonymous):

its fine lol u can just look at the graph

OpenStudy (anonymous):

i didn't even see the graph lol how salty am i lol

OpenStudy (anonymous):

lol i do that sometimes 2 .. Its fine:)